OpenStudy (anonymous):

What is the equation of the line that contains the points (3, –6) and (–3, 0)?

OpenStudy (anonymous):

\[ \frac{y-(-6)}{x-(3)} = \frac{(0)-(-6)}{(-3)-(3)} \]

OpenStudy (anonymous):

So would that also mean y=x-3?

OpenStudy (anonymous):

dunno, I didn't simplify it

OpenStudy (triciaal):

no check the left side